OpenStudy (lukecrayonz):

((y^2-y-12)/(y^2-6y+8))/((y^2+2y-15)/(y^2-7y +10))

OpenStudy (anonymous):

I can't read your equation

OpenStudy (lukecrayonz):

http://screensnapr.com/v/IMcvqY.png

OpenStudy (anonymous):

\[\frac{\left(y^2-7 y+10\right) \left(y^2-y-12\right)}{\left(y^2-6 y+8\right) \left(y^2+2 y-15\right)}=\frac{(y-5) (y+3)}{(y-3) (y+5)} \]

jhonyy9 (jhonyy9):

(y-4)(y+3) (y-5)(y-2) ---------- * --------- = ((y-5)(y+3))/((y+5)(y-3)) (y-4)(y-2) (y+5)(y-3)